OpenStudy (anonymous):

What is the value of x? -15 - x = -6 i think its 9 am i wrong ?

OpenStudy (anonymous):

Check your solution by substituting it in for x and seeing if it comes out true. -9 you think? Let's see. -15 - (-9) = -6 -15 + 9 = -6 -6 = -6 True!! So yes, your solution x = 9 is correct.

OpenStudy (anonymous):

alright thanks man @Cosmichaotic